OpenStudy (anonymous):

pleasee helppppp!!!!!!!!!!!!!!! Find all solutions to the equation. cos x = sin x

OpenStudy (mertsj):

Divide both sides by cos x.

OpenStudy (anonymous):

or look on the unit circle and see where the first coordinate is equal to the second coordinate

geerky42 (geerky42):

\[x = \dfrac{\pi}{4}+k\pi, k \in \mathbb{Z}\]

OpenStudy (anonymous):

so would it be like \[\pi/4 +n \pi | n= +/- 1, +/- 2....\]

geerky42 (geerky42):

Yeah.

OpenStudy (anonymous):

okay thank you :))